It costs a pretzel vendor $20 per day to rent a cart and $0.75 for each pretzel.
(A) Find the cost of selling x pretzels in 1 day.
(B) What is the cost of selling 150 pretzels in 1 day?
(C) If the daily cost is $275, how many pretzels were sold that day?

Answer by Theo(13342)

the formula for selling x pretzels in a day is:

y = 20 + .75 * x

if the number of pretzels sold is 150, then the formula becomes:

y = 20 + .75 * 150

solve for y to get:

y = 20 + 112.5 = 132.5

if the daily cost is 275, then the number of pretzels sold is calculated below.

start with 275 = 20 + .75 * x
subtract 275 from both sides of the equation to get:
255 = .75 * x
divide both sides by .75 to get:
340 = x

340 pretzels are sold if the total cost for the day is 275.

y = 20 + .75 * 340 = 275, confirming the solution is correct.